5 * 3 * 1
* 4 * 2
5 * 3
* 2
5
1
Expert's answer
2013-03-12T09:48:46-0400
(5 * 3 * 1)*( 4 * 2*5) *( 3* 2*5)=15*40*30=15*1200=(15*12)*100=180*100=18000